# backgroundcolor=gray!3, roundcorner=3pt

### v1

@@ 4.4 Transforming fused policies from the certainty domain to the probability domain
Calculating probabilities. We take the fused policy in Tab. 6 and for each tuple, we apply Equation 13. Tab. 7 shows the result of this transformation and the new policy that considers advice [1,1$\rightarrow$-2. Consider, e.g., the Down action in state [0, 1]. Here, $\omega=(0.143,0.857,0.0,0.25))$, and from Equation 13, $p=b+au=0.143+0.25\times 0.0=0.143$ follows.
As seen by the highlighted values, the advice affects the actions of neighboring states that lead to state [1,1], i.e., [0,1] - Down, [1,0] - Right, [1,2] - Left, [2,1] - Up. Tab. 7(a) and Tab. 7(b) show the policy before and after normalization, respectively.
Normalization. The resulting probabilities in row [0,1] violate invariant Equation 14 as $\sum_{a\in A}p(a|s=[0,1])=0.893$. Thus, we normalize by Equation 15 and multiply each probability by $\frac{1}{\sum_{a\in A}p(a|s=[0,1])}$, resulting in the probability vector (0.28, 0.16, 0.28, 0.28), as shown in the [0,1] row of Tab. 7(b).
Full example. To round out the running example, we show the full policy after considering every piece of advice in Listing 2. Tab. 8 shows the the new probabilities before (Tab. 8(a)) and after (Tab. 8(b)) normalization. As seen, some advice does not change the probabilities of the policy much, e.g., in the [0,3] - Down row of Tab. 8(a) (0.217, previously 0.250); while other pieces of advice have substantial impact, e.g., in [3,2] - Right (0348, previously 0.250). This is due to the uncertainty of the advice we modeled by the distance between the advisor and the advised state. In the running example, the advisor was situated in the bottom-left corner, and therefore, as a consequence, their advice about the distant [1,3] cell impacted the Down action of cell [0,3] less than the advice about the closer [3,3] cell that impacted the Right action of cell [3,2].
Table 7: Fused policy of the Frozen Lake running example in the probability domain after advice [1,1$\rightarrow$-2 affecting the actions of neighboring states [0,1], [1,0], [1,2], [2,1] that lead to state [1,1]
@@
Table 8: Complete fused policy of the Frozen Lake running example
@@
Visualization of results. The final policy in Tab. 8(b) is visualized in Fig. 5, with the major changes highlighted. As seen, the negative advice about the hole in [1,1] resulted in a strong decrease in the corresponding probabilities of actions from neighboring cells, while the negative advice about the hole in [1,3] resulted in slightly smaller decrease due to higher uncertainty of the advice. Conversely, the positive advice about the goal in [3,3] coupled with high certainty resulted in substantial positive bias that will guide the agent around the hole to the goal. Note that this policy is only the input to the agent’s learning phase, and the preferences in each state have yet to be reinforced. For example, in state [0,2], there is no change based on the advice, but surely, the agent will learn not to move down toward a hole but rather, to the right.
@@ 5.1.1 Manual experimentation
To experiment with various map configurations in a reproducible way, we developed a map generator that is able to generate maps of arbitrary size and hole density and disperse holes on the map randomly, controlled by a user-defined seed.
Eventually, we choose a $12\times 12$ map with 20% hole density (which is the default density in OpenAI Gymnasium) under seed 63. The final map is shown in Appendix A. On this map, we observe that the behavior of the unadvised RL agent is appropriate after 5 00050005\,0005 000 episodes. Thus, we set 5 00050005\,0005 000 as the minimum number of episodes we use in our evaluation.
@@ 5.1.2 Scenario
To generate such a large map, we developed a map generator in Python that allows for generating maps of arbitrary size and hole ratio. Holes are distributed randomly on the lake. For replicability, the generator can be parameterized with a seed. The generator outputs an Excel file for human inspection and the standard graphical rendering of the generated environment. The Excel file is then used as input for each experiment execution.
@@ 5.1.3 Experimental setup
Advice strategy Once, at the beginning of the experiment
@@
Agent type {Random, Non-adviced, Advised}
Source of advice {Synthetic (oracle), Human}
Synthetic advice type {Every known cell (“ALL”), Just the holes and goal (“HOLES&GOAL”)}
Human advice quota {5%, 10%}
Degree of uncertainty ($u$) {$0.2k~{}|~{}k\in{0..5}$}
Fixed parameters. In this work, we choose discrete policy gradient as the reinforcement learning method. Policy-based methods represent the policy explicitly, which allows for directly investigating the effects of advice on the policy.
@@
Variables. We execute experiments for each agent type to compare their performance. The random agent randomly samples from the available actions. Non-advised agents maintain a policy. Advised agents maintain a policy, and in the first step, they fuse the advice into their initial policy. In each case, the initial policy is the one in which the probability of choosing an action is uniformly distributed.
We execute experiments with different sources of advice. In the synthetic mode, a hypothetical and idealized oracle formulates its advice. We use two advice types in the synthetic mode: advice about every cell (“ALL”); and advice about the holes and the goal (“HOLES&GOAL”). In the human advice mode, a human formulates realistic advice based on partial knowledge. We tie the number of advices to the size of the problem and use 5% and 10% of the number of cells as the advice quota.
Both in synthetic and human advice mode, we execute experiments with different degree of uncertainty ($u$) that sweeps through the [0.0, 1.0] interval in 0.2 increments. Here, $u=0$ reduces to classic probability, and $u=1$ represents a uniform distribution [28].
@@
We execute experiments using each combination of parameters defined above. We run each experiment 30 times to achieve sufficient statistical power. The results of experiments are saved as csv files.
@@
5.1.4 Threats to validity and limitations
Construct validity
Observations of improved exploration performance may be artifacts of fortunately aligned problems (here: maps) or a few fortunate pieces of advice rather than artifacts of systematic effects. To mitigate this threat to validity, first, we experimented with different map layouts and observed the same performance trends. To facilitate consistent manual experimentation, we developed a map generator that generates map layouts in a reproducible way. Second, we ran our scaled-up experiments with a variety of settings, including both human and synthetic advisors.
Internal validity
Evaluating the approach with human advisors inevitably poses threats to internal validity. Specifically, maturation and testing might have influenced our results as the authors experimented with advising strategies and advice values. For example, cells that are not holes or goals are labeled freely with -1, 0, or 1, and a human might learn how to misuse these values to achieve a specific behavior by the agent. We mitigated these threats by defining clear advise score rules, which are also used by the synthetic advisor.
External validity
We chose only one flavor of reinforcement learning—a policy-based method—to demonstrate our approach, and we did not provide evidence that our approach can be applied to other reinforcement learning methods. However, we are reasonably confident that our methods translate to reinforcement learning methods where the policy is not explicitly represented. For example, in value-based methods, advice can be fused into the expected cumulative rewards in a given state, similar to the approach presented in this article. Nonetheless, in future work, we will validate this assumption.
Limitations
We used a relatively simple advising strategy, in which advice is provided before the exploration. It is reasonable to assume that more interactive advising modes might shed light on additional benefits and challenges in guided reinforcement learning. However, we are reasonably confident that the scope of this article was sufficient to draw key takeaways that hold in more interactive reinforcement learning setups as well. Nonetheless, we will experiment with interactive advising modes in future work.
@@ 5.2 Results
In the following, we report the results of our experiments. First, we have a look at the performance of synthetic advisor (Sec. 5.2.1), and subsequently, we review the performance of real human advice (Sec. 5.2.2). In both cases, we first investigate the performance in terms of the cumulative reward, and then we look at the final policy.
As the key takeaway of our work, Tab. 10 highlights that even uncertain advice at $u=0.6$ improves the performance of the agent. We observe mixed results with very uncertain advice at $u=0.8$; here, the volume of advice by the synthetic agent somewhat compensates for the low certainty, and keeps the performance slightly above that of the unadvised agent, but a human advisor with limited advice quota and such high uncertainty fails to improve over the unadvised agent. Thus, $u=0.6$ seems to be the level of uncertainty at which opinion-guided agents consistently outperform the unadvised agent.
Table 10: Cumulative rewards by experiment. (Bold is best at the given level of certainty. Red is below the performance of the unadvised agent.)
Synthetic Human
All Holes 10% 5%
Advice@u=0.0 8 778.033 9 045.933 8 541.633 6 674.333
Advice@u=0.2 7 584.600 6 904.033 6 226.367 3 527.967
Advice@u=0.4 4 702.267 3 023.000 4 345.033 1 539.600
Advice@u=0.6 3 102.833 1 789.500 2 373.800 855.933
Advice@u=0.8 1 638.000 909.967 285.400 527.867
Advice@u=1.0 489.600 183.833 351.833 461.633
Random 0.100
No advice 823.367
5.2.1 Performance of guidance with synthetic advisor
@@
Fig. 6 shows the cumulative rewards of the agent with a synthetic advisor under two different advice strategies. Each plot shows the random agent and the agent without advice, compared to advised agents at five levels of uncertainty ($u=0.0,0.2,0.4,0.6,0.8,1.0$).
@@
(a) Advice strategy: ALL
(b) Advice strategy: HOLES&GOAL
@@
(c) Advice strategy: ALL
(d) Advice strategy: HOLES&GOAL
Figure 6: Cumulative reward with a synthetic advisor
Clear performance improvement in advised agents. As the most important takeaways, we observe that (i) advised agents clearly outperform the unadvised agent and that (ii) performance improves with the level of certainty of the advice. This can be seen from the lines corresponding to lower uncertainty consistently being over lines corresponding to higher uncertainty. As expected, the random agent hardly accumulates any reward as it follows a random trajectory without learning.
@@
Even substantially uncertain advice makes a difference. Particularly demonstrated in the $u=0.6$ case, agents with low-certainty advice can still outperform the unadvised agent. In very uncertain cases, i.e., $u=0.8$ and $u=1.0$, the performance of the advised agent is similar to that of the unadvised agent. The extensive “all” advise strategy compensates for low certainty and renders the advised agent in the $u=0.8$ case more performant than the unadvised agent.
Learning dynamics: more rapid learning. The log scale charts further show the two main phases of the learning process. In the first phase, approximately between episodes 0–2 000, we see a steep increase, indicating fast learning. In the second phase, approximately after episode 2 000, the slopes become more gradual, indicating a slowed-down collection of reward and a likely convergence to the maximum potential performance. Performance differences remain consistent throughout the experiment. The effect of advice, thus, is shifting the exponentially growing initial phase earlier in the learning process.
Advice strategies: difference shows only in low-certainty cases. There is not much difference between the two advice strategies in high-certainty cases, as shown by the very similar green ($u=0.0$) and red ($u=0.2$) lines in Fig. 5(a) and Fig. 5(b). In more uncertain cases ($u>0.2$), there is a noticeable performance deterioration in the holes-and-goal strategy, as neither line reaches the same eventual cumulative reward as in the “all” strategy.
@@
Advice strategy: HOLES&GOAL
@@
Figure 7: Heatmap visualization of the final learned policy with a synthetic advisor
@@
Better reinforced policy. One of the two key improvements is indicated by the darker shades of cells compared to the unadvised agent. Darker shades correspond to higher probabilities of the indicated action in a cell and an overall better reinforced beneficial policy. By providing advice, the agent navigated through the state space more efficiently and had time (episodes) to explore crucial cells, e.g., the ones around holes and the goal. The “holes&goal” strategy reinforces some key cells better than the extensive “all” strategy, e.g., around the goal and around the starting point.
More thorough exploration. The other key improvement is shown by the higher number of cells explored in advised strategies as both the “all” and “holes&goal” advice strategies cover more cells compared to the unadvised agent. This is possible due to the advice driving the agent in crucial situations, resulting in fewer failed attempts. The “holes&goal” strategy allows to explore more states at higher levels of certainty than the extensive “all” strategy (cf. Fig. 7(e) and Fig. 7(c)); however, this relation turns around in the lower-certainty case (cf. Fig. 7(b) and Fig. 7(d)).
Effects clear even at relatively low certainty. At $u=0.4$, we still see demonstrated improvements compared to the unadvised agent. That is, even fairly uncertain advice has a clear positive impact on the policy.
5.2.2 Performance of guidance with human advisor

### v2

@@ 4.4 Transforming opinion-infused policies from the certainty domain to the probability domain
Calculating probabilities
We take the fusion of the advice and the policy in Tab. 6 and for each tuple, we apply Equation 16. Tab. 7 shows the result of this transformation and the new policy that considers advice [1,1$\rightarrow$-2. Consider, e.g., the Down action in state [0,1]. Here, $\omega=(0.143,0.857,0.0,0.25))$, and from Equation 16, $p=b+au=0.143+0.25\times 0.0=0.143$ follows.
As seen by the highlighted values, the advice affects the actions of neighboring states that lead to state [1,1], i.e., [0,1]Down, [1,0]Right, [1,2]Left, [2,1]Up. Tab. 7(a) and Tab. 7(b) show the policy before and after normalization, respectively.
Normalization
The resulting probabilities in row [0,1] violate invariant Equation 17 as $\sum_{a\in A}p(a|s=[0,1])=0.893$. Thus, we normalize by Equation 18 and multiply each probability by $\frac{1}{\sum_{a\in A}p(a|s=[0,1])}$, resulting in the probability vector (0.28, 0.16, 0.28, 0.28), as shown in the [0,1] row of Tab. 7(b).
Full example
To round out the running example, we show the full policy after considering every piece of advice in Listing 2. Tab. 8 shows the the new probabilities before (Tab. 8(a)) and after (Tab. 8(b)) normalization. As seen, some advice does not change the probabilities of the policy much, e.g., in the [0,3]Down row of Tab. 8(a) (0.217, previously 0.250); while other pieces of advice have substantial impact, e.g., in [3,2]Right (0348, previously 0.250). This is due to the uncertainty of the advice we modeled by the distance between the advisor and the advised state. In the running example, the advisor was situated in the bottom-left corner, and therefore, as a consequence, their advice about the distant [1,3] cell impacted the Down action of cell [0,3] less than the advice about the closer [3,3] cell that impacted the Right action of cell [3,2].
Table 7: Shaped policy of the Frozen Lake running example in the probability domain after advice [1,1$\rightarrow$-2 affecting the actions of neighboring states [0,1], [1,0], [1,2], [2,1] that lead to state [1,1]
@@
Table 8: Shaped policy of the Frozen Lake running example
@@
Visualization of results
The final policy in Tab. 8(b) is visualized in Fig. 5, with the major changes highlighted. As seen, the negative advice about the hole in [1,1] resulted in a strong decrease in the corresponding probabilities of actions from neighboring cells, while the negative advice about the hole in [1,3] resulted in slightly smaller decrease due to higher uncertainty of the advice. Conversely, the positive advice about the goal in [3,3] coupled with high certainty resulted in substantial positive bias that will guide the agent around the hole to the goal. Note that this policy is only the input to the agent’s learning phase, and the preferences in each state have yet to be reinforced. For example, in state [0,2], there is no change based on the advice, but surely, the agent will learn not to move down toward a hole but rather, to the right.
Reflection
The agent’s policy is affected by the advice. Positive advice by opinions with belief being greater than base rate ($b>a$) causes the shaped policy to bias towards more belief to take actions that lead to the state that has been advised as a beneficial one. This bias manifests in an increased probability of the agent taking corresponding actions that lead to the advised state from its neighboring states. Such cases can be observed around the goal in the running example. In contrast, negative advice ($b<a$) shapes the policy in a way that it drives the agent way from advised states by decreasing the probability of the agent taking corresponding actions that to the advised state from its neighboring states. Such cases can be observed around the holes in the running example. The level of uncertainty has a profound impact on the shaped policy, as demonstrated by the difference between the weight of negative advice about state [1,1] and [1,3].
@@ 5.1.1 Manual experimentation
To experiment with various map configurations in a reproducible way, we developed a map generator that is able to generate maps of arbitrary size and hole density and disperse holes on the map randomly, controlled by a user-defined seed. The map generator is available in the replication package.
Eventually, we choose a $12\times 12$ map with 20% hole density (which is the default density in OpenAI Gymnasium) under seed 63. The final map is shown in Appendix A. On this map, we observe that the behavior of the unadvised RL agent is appropriate after 5 00050005\,0005 000 episodes. Thus, we will choose a sufficiently higher number of episodes in our experiments, as explained below.
@@ 5.1.2 Scenario
To generate such a large map, we developed a map generator in Python that allows for generating maps of arbitrary size and hole ratio. Holes are distributed randomly on the lake. For replicability, the generator can be parameterized with a seed. The generator outputs an Excel file for human inspection and the standard graphical rendering of the generated environment. The Excel file is then used as input for each experiment execution. The map generator is available in the replication package.
@@ 5.1.3 Experimental setup
Learning rate 0.9
Discount factor 1.0
Advice strategy Once at the beginning of the experiment
@@
Agent type {Random, Unadvised, Advised}
Source of advice {Oracle, Single human, Cooperating humans}
Advice quota – oracle {100(“All”), 20% (“Holes&Goal”)}
Advice quota – single human {10%, 5%}
Advice quota – coop. human {10each, 5% each}
Uncertainty (oracle and single human) {$0.2k~{}|~{}k\in{0..4}$}
Uncertainty (cooperative humans) Calculated dynamically from a distance metric
Cooperative advice type {Sequential cooperation, parallel cooperation}
Fixed parameters
In this work, we choose discrete policy gradient as the reinforcement learning method. Policy-based methods represent the policy explicitly, which allows for directly investigating the effects of advice on the policy. To mitigate threats to validity, we use meaningful hyper parameters of the RL algorithm in order to allow the unadvised agent to perform as well as possible. Specifically, we set the learning rate to 0.9, and the discount factor to 1.0.
@@
Finally, we fix the problem size to a $12\times 12$, as explained in Sec. 5.1.1.
Variables
We execute experiments for each agent type to compare their performance. The random agent randomly samples from the available actions. Unadvised agents maintain a policy. Advised agents maintain a policy, and before their first training episode, advice is fused into their initial policy. In each case, the initial policy is the one in which the probability of choosing an action is uniformly distributed.
We conduct experiments with different sources of advice. To form a ground truth, we first experiment with an idealized oracle with full information about the problem space. The oracle provides synthetic advice based on pre-defined rules as follows. If the cell is a goal, its value is +2. If the cell if a hole, the value is -2. If the cell has no neighboring holes, its value is +1. If the cell has one neighboring hole, its value is 0. If the cell has more than one neighboring hole, its value is -1. In the single human mode, a human advisor provides advice, and we modulate the uncertainty of their advice, similar to the oracle’s advice. In the cooperative mode, advice is provided by two human advisors who have partial and complementary information about the problem. Thus, this experimental mode is highly realistic. The human advisor labels hole and goal cells similarly to the oracle, but for other cells, may use values ±$\pm 1,0$ differently than the oracle.
We use different advice quotas for different advisor types. We evaluate the oracle by two advice quotas: first, by providing the agent with advice about all the cells, i.e., 100% quota; second, by providing the agent with advice about the holes and the goal on the map, which amounts to about 20% of cells (see Sec. 5.1.2). We evaluate the single human advisor by two advice quotas: first, by providing the agent with advice about 10% of the cells; second, by providing the agent with advice about 5% of the cells. We evaluate the cooperating human advisors by two human advice quotas: first, by allowing each advisor 10% advice quota (thus, 20% in total); second, by allowing each advisor 5% advice quota (thus, 10% in total).
We synthetically modulate the degree of uncertainty of the advice of the oracle and the single human by sweeping through the [0.0, 1.0) interval in 0.2 increments. Here, $u=0.0$ reduces to classic probability, and $u=1$ represents a uniform distribution [33]. This choice allows us to directly compare the oracle with the single human advisor at different levels of uncertainty.
In the cooperative mode, uncertainty is calculated from the distance between the advisor and the location of the advice, in accordance with Sec. 4.2.1. We use the two dimensional Manhattan distance, defined as follows.
$D((x_{1},y_{1}),(x_{2},y_{2}))=|x_{1}-x_{2}|+|y_{1}-y_{2}|.$
Here, $(x_{1},y_{1})$ is the location of the advisor. We assume a linear certainty discount function with $\tau=1$, as explained in Equation 5. That is, uncertainty of an advisor in one of the corners of the grid reaches its maximum in the opposing corner. Thus, uncertainty at any point $(x_{2},y_{2})$ is defined as follows.
$u(x_{2},y_{2})=D((x_{1},y_{1}),(x_{2},y_{2}))\cdot\frac{1}{D((x_{1},y_{1}),(x_{n},y_{n}))}.$
Finally, we define two types of cooperative advice with partial information. In sequential cooperation, one human advises the agent in the first half of its mission, and subsequently, the other human advises the agent in the second half of its mission. To achieve this, we place the first human advisor in the top-left corner of the grid (start), and the second human advisor in the bottom-right corner of the grid (goal). As the agent follows the main diagonal of the grid world, the first advisor’s input will gradually lose its influence due to the distance explained above; and the second advisor’s input will gradually gain more influence. In parallel collaboration, both humans provide advice throughout the agent’s entire mission. To achieve this, we place one human advisor in the top-right corner of the grid, and the other human advisor in the bottom-left corner of the grid. By that, both advisors are able to provide advice throughout the entire start-goal trajectory of the agent. However, due to the distance metric defined above, both advisors will focus on their half of the grid world, i.e., the triangle under and above the diagonal, respectively. Clearly, this advice mode is the most realistic one of the three as in real settings, uncertainty pertains to individual pieces of advice, and cooperation among stakeholders and experts is the usual way to solve complex problems.
@@
We execute experiments using each combination of parameters defined above. We run each experiment 30 times to achieve sufficient statistical power. The results of experiments are saved as csv files. These data files are available in the replication package.
@@ 5.2 Results
In the following, we report the results of our experiments by reviewing the performance of guidance by the oracle (Sec. 5.2.1), single human advisor (Sec. 5.2.2), and cooperative human advisors (Sec. 5.2.3). In each case, we first investigate the performance in terms of the cumulative reward, and then we look at the final policy.
Main takeaways
Tab. 10 and Fig. 6 report the cumulative rewards in different evaluation modes and highlights the key takeaways of our work.
We note that every advised agent performs better than the unadvised agent (607.267 mean cumulative reward), even at high levels of uncertainty.
We observe comparable results between the fully informed (100% quota) oracle and the 10% quota single human case at low levels of uncertainty. In fact, at $u=0.2$, the human advice with 10% quota outperforms the oracle, even with 100% quota (8 607.500 vs 8 511.367).
Finally, to compare cooperative performance with other experiments, we show the results of the four realistic cooperative experiments aligned with the comparable oracle or single human experiments with the same advice quota. For example, sequential cooperation with 2$\times$10quota is comparable to the oracle with 20% quota. In this comparison, performances align best at $u=0.2$, where cooperating humans with partial information perform nearly as well as the oracle with complete information about the problem space (7 551.567 vs 7 835.267). In general, sequential cooperation outperforms parallel cooperation; and parallel cooperation at 5% outperforms the fully-informed single human advisor at 10%.
Table 10: Cumulative rewards by experiment. Bold is best at the given level of certainty. Unadvised = 607.267. Random = 0.100. Cooperative modes aligned with the best matching uncertainty level of the comparable oracle or single human mode.
Oracle Single human Coop. – Sequential Coop. – Parallel
u 100% (All) 20% (H&G) 10% 5% 10% 5% 10% 5%
0.0 9 386.467 9 443.733 8 907.733 7 576.067
0.2 8 511.367 7 835.267 8 607.500 4 656.633 7 551.567
0.4 7 476.633 4 218.000 4 727.433 1 718.300 5 544.100 4 924.967
0.6 2 751.833 2 089.433 2 737.600 1 360.267 2 867.633
0.8 1 454.400 878.467 1 907.933 629.367
In the following, we elaborate on the experimental results in detail.
Linear scale
(a) Oracle 100%
(b) Oracle 20%
(c) Human 10%
(d) Human 5%
(e) Coop $10\%$
(f) Coop $5\%$
Log scale
(g) Oracle 100%
(h) Oracle 20%
(i) Human 10%
(j) Human 5%
(k) Coop: $10\%$
(l) Coop $5\%$
Figure 6: Performance comparison of every experiment in terms of cumulative reward
5.2.1 Performance of guidance by an oracle
Advisor type
In this experiment, we use an oracle, i.e., synthetic advice that assigns values to states based on pre-defined rules (as explained in Sec. 5.1.3).
Advice quota
We evaluate two advice quotas: 100% (the agent is provided with advice about all the cells by an idealized oracle) and 20% (the agent is provided with advice about the terminating states, i.e., the holes and goal).
Uncertainty calibration
We execute each experiment at multiple characteristic levels of uncertainty, sweeping through the {0.0, 0.2, 0.4, 0.6, 0.8} range. The case of complete uncertainty $u=1.0$ is not tested as these opinions carry no information that can be used by the agent. [^3] It can also be shown that in the fusion operator of choice in this work (BCF), $u=1.0$ results in an idempotent transformation, which follows trivially from Equations 1015. The proof is left as an exercise for the reader.
@@ 5.2.1 Performance of guidance by an oracle
Fig. 7 shows the cumulative rewards of the agent with synthetic advice from the oracle under the two advice quotas.
@@
(a) Advice quota: 100% (All)
(b) Advice quota: 20% (Holes&Goal)
@@
(c) Advice quota: 100% (All)
(d) Advice quota: 20% (Holes&Goal)
Figure 7: Cumulative reward with an oracle as the advisor
Clear performance improvement in advised agents
As the most important takeaways, we observe that (i) advised agents clearly outperform the unadvised agent and that (ii) performance improves with the level of certainty of the advice. This can be seen from the lines corresponding to lower uncertainty consistently being over lines corresponding to higher uncertainty. As expected, the random agent hardly accumulates any reward as it follows a random trajectory without learning.
@@
Even substantially uncertain advice makes a difference
Demonstrated in moderate ($u=0.6$) and high ($u=0.8$) uncertainty cases, advised agents still outperform the unadvised agent. The 100% advice quota compensates for low certainty and renders the advised agent in the $u=0.8$ case more performant than the unadvised agent. This advantage is lost at 20% quota as shown in Fig. 6(b) by the similarity of the 0.8 and “No advice” lines.
Learning dynamics: more rapid learning
The log scale charts further show the two main phases of the learning process. In the first phase, approximately between episodes 0–2 000, we see a steep increase, indicating fast learning. The learning phase is substantially steeper as uncertainty decreases. In the second phase, approximately after episode 2 000, the slopes become more gradual, indicating a slowed-down collection of reward and a likely convergence to the maximum potential performance. Performance differences remain consistent throughout the experiment. The effect of advice, thus, is shifting the exponentially growing initial phase earlier in the learning process.
Advice quotas: difference shows only in low-certainty cases
There is not much difference between the two advice quotas in high-certainty cases, as shown by the very similar lines of $u=0.0$ and $u=0.2$ in Fig. 6(a) and Fig. 6(b). In fact, as shown in Tab. 10, the oracle with 20% quota slightly outperforms the oracle with 100% quota at $u=0.0$. (+0.61%) The most noticeable performance deterioration is observed at $u=0.4$, when performance drops by 43.6% between the 100% quota and 20% quota.
@@
Advice quota: 20% (HOLES&GOAL)
@@
Figure 8: Heatmap visualization of the final learned policy with an oracle as the advisor
@@
Better reinforced policy
One of the two key improvements is indicated by the darker shades of cells compared to the unadvised agent. Darker shades correspond to higher probabilities of the indicated action in a cell and an overall better reinforced beneficial policy. By providing advice, the agent navigated through the state space more efficiently and had time (episodes) to explore crucial cells, e.g., the ones around holes and the goal. The agent with 20% advice quota reinforces some key decisions better than the agent with 100% advice quota, e.g., around the goal and around the starting point.
More thorough exploration
The other key improvement is shown by the higher number of cells explored by advised agents, as both advice quotas result in more cells being covered, compared to the unadvised agent. The 20% advice quota allows to explore more states at higher levels of certainty than the 100% advice quota (cf. Fig. 8(e) and Fig. 8(c)); however, this relation turns around in the lower-certainty case (cf. Fig. 8(b) and Fig. 8(d)).
Effects clear even at relatively low certainty
At $u=0.4$, we still see demonstrated improvements compared to the unadvised agent. That is, even fairly uncertain advice has a clear positive impact on the policy.
5.2.2 Performance of guidance by a single human advisor
Advisor type
In this experiment, we use the advice of a single human advisor.
Advice quota
We evaluate two advice quotas. In the first case, the human provides advice about 10% of the cells at their discretion. In the second case, the human provides advice about 5% of the cells at their discretion.
Uncertainty calibration
We execute each experiment at multiple characteristic levels of uncertainty, sweeping through the {0.0, 0.2, 0.4, 0.6, 0.8} range. The case of complete uncertainty $u=1.0$ is not tested as these opinions carry no information that can be used by the agent.